OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 288 cu. in. what i am doing is multiplying 288 by 4/3 to get back to the origional radius squared i get 384 sence that is alredy the radius squared i just plug it into the sphere area formula (4 pi radius squared) and multiply it by 4 and i get 1536 pi i know that is incorrect because my answer choices are 48 pi 144 pi 20,736pi what am i doing wrong?

OpenStudy (anonymous):

The key idea is that both the surface area and the volume are related to the radius

OpenStudy (anonymous):

so solve for the radius and then find the surface area

OpenStudy (anonymous):

Volume = (4/3)pi r^3 = 288

OpenStudy (anonymous):

r^3 = 216/pi

OpenStudy (anonymous):

r = cuberoot(216/pi) = cuberoot(216)/cuberoot(pi) = 6/cuberoot(pi)

OpenStudy (anonymous):

Surface area = 4 pi r^2 = 4pi(6/cuberoot(pi))^2 = 144pi/cuberoot(pi^2) = 144cuberoot(pi)

OpenStudy (anonymous):

Answer=144cuberoot(pi)

OpenStudy (anonymous):

Have you got it ?

OpenStudy (anonymous):

so next time i should solve 4 the radius and not just plug it in

OpenStudy (anonymous):

ok thanks a million!!:)